What is magnitude? Give me full detail about this.


 

Grade:12th Pass

4 Answers

Karan Malhotra
11 years ago

Magnitude is defined as the value of a physical quantity.For example if a body is travelling with a speed of 5m/s,then the magnitude of the speed is 5.Magnitude is often represented like this-|Q|,where Q isa physical quantity.For example |speed|=5.

MUKESH chinna YADAV
11 years ago

magnitude is the size or lenth of the particle

vikash kumar
11 years ago

our teacher say that magnitude is not size or length.He want to talk with you.Please give me your mobile number or full detail about magnitude, that satisfy our teacher.

G Ashwin -
11 years ago

Magnitude refers to the real no. value.

eg. magnitude of 3i+4j+2k = root (3^2+4^2+2^2)

                                       = root(29)

                                       =  5.38